OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPPPPPP Which point lies on the locus of points equidistant from A (2,0) and B (6,0) ?

OpenStudy (anonymous):

this means that \[(x-2)^2+y^2=(x-6)^2+y^2\]

OpenStudy (anonymous):

\[x^2-4x+4+y^2=x^2-12x+36+y^2\] \[-4x+4=-12x+36\] \[8x=32\] \[x=4\]

OpenStudy (anonymous):

it is the vertical line \(x=4\)

OpenStudy (anonymous):

So.. would it be (0,4) (2,5) or (4,5) ?

OpenStudy (anonymous):

so \((4,5)\) should work

OpenStudy (anonymous):

any point where \(x=4\)